Alkalinity: The approximate alkalinity of a water sample is 2.21 x 10-2 M at pH=10.33. Determine the concentration of carbonate (CO32-) in M and mg/L as CaCO3. (pKa=10.33).

Answer: [CO32-] = 7.37x10-3 M and 737 mg/L as CaCO3

Please show the solution to this problem.
Answer #1

Given,

The alkalinity of a water sample = 2.21 x 10-2 M at pH=10.33.

Alkalinity of water is the ratio between ionic weight of CO32- ions and the molecular mass of CaCO3 .

Alkalinity = (concentration of CO32- x molecular mass of CaCO3 ) / Ionic weight of CO32-

We know,

molecular mass of CaCO3 = 100 g/mol

Ionic weight of CO32- = 60 g/mol

Substituting the values,

2.21 x 10-2 = ( c x 100 ) / 60

c =( 2.21 x 10-2 x 60 ) / 100 = 0.01326 M = 1.326 mg/L

I hope the provided answer may be wrong. I suggest you to recheck it.

Best Wishes.